OpenStudy (anonymous):

(3)(-7)(0)(-9)= Find the product

OpenStudy (anonymous):

0. Anything times zero is zero. 3*-7*0 is 0; times -9 is zero. -9*-3*0*3 is also 0

OpenStudy (anonymous):

*(-9)(3)(0)(-7)

OpenStudy (anonymous):

What type of number results when a negative number is multiplied by a positive number and vice versa? A. A negative number B. Zero C. A positive number

OpenStudy (anonymous):

a negative times a positive is a negative. a negative times a negative is a positive.

OpenStudy (anonymous):

-5*2 =-10 -5*-2=10

OpenStudy (anonymous):

Find the Product [6+(-8)]x(-5)=

OpenStudy (anonymous):

(6+-8)x *-5 What is 6-8? Add an x to that number and multiply by -5

OpenStudy (anonymous):

6+-8 is 6-8

OpenStudy (anonymous):

Solve. A hockey player's plus-minus rating compares the number of goals scored by his or her team (plus) to the number scored by the opponents (minus) while he or she is playing. In 1996, eight players on the San Jose Sharks had an average plus-minus rating of -17. What was the total of their ratings?